A telephoto lens with focal length of 120mm has f-numbers ranging from f/2.8 to f/22.

What is the diameter of the lens aperture at f/2.8?

What is the diameter of the lens aperture at f/22?
Answer #1

f-number = focal length/lens diameter.


so: focal length/f--number = lens diameter


(120 mm)/2.8 = 42.86 mm


(150 mm)/22 = 5.45mm
Answer #2

f-number = focal length/lens diameter.

so: focal length/f--number = lens diameter

(120 mm)/2.8 = 42.85 mm

(120 mm)/22 =5.454 mm